Evaluate the following indefinite integral. | asin (a)cos(2) de Do not include + in your answer. For example, if you found
Answer #1


[4sinº (x)cos(x) dx = 4((sin (x))° (cos(x) dx) Let, sin x=u cos xdx = du = 4/uºdu+C 7 +C 4 = -u +C [ 4 sinº ( x) cos(x)dx= s